# Introducing a Physics-informed Deep Learning Framework for Bridge Scour Prediction

### v2

@@ 2.3.4 Long Short-Term Memory Networks (LSTM)
In this study, we implement a one-layer LSTM model with 128 hidden units as shown in Fig. 5. The input sequence is processed by the LSTM layer and then passed through a dense (fully connected) layer with a linear activation function and eventually reshaped into the output sequence.
Figure 5: The architecture of LSTM model. An LSTM Memory Unit below showcasing the input ($x_{t}$), hidden ($h_{t}$), and cell state ($c_{t}$) vectors, with gates (forget, input, and output) that dynamically regulate the flow of information through the unit to maintain and update the memory.
2.3.5 Convolutional Neural Network (CNN)
@@
In this study, we develop a CNN model with two 2D convolutional layers to show the adaptability of CNNs for processing multidimensional time series data. The two CNN layers have a kernel size of 5, padding size of 2, and output sizes of 128 and 256, respectively. Also, batch normalization layers are added after these layers, and one following dense layer is responsible for producing the final prediction.
Figure 6: The architecture of CNN model.
2.3.6 NLinear
The NLinear neural network is a variant within the Long-term Time Series Forecasting Linear (LTSF-Linear) framework proposed by Zeng et al. (2023), which uses only one dense layer to model multi-step time series forecasting problems. Notably, this variant is designed to improve the performance in the presence of distributional shifts of temporal data. This phenomenon refers to the temporal fluctuations in data properties caused by various factors, such as seasonal patterns and short-term trends, which reduces the generalization ability of predictive models trained on past observations. The NLinear model (see Fig. 7) mitigates distribution shifts by centring the input sequence around its last value, both before and after the linear transformation. It subtracts the last value from the sequence, processes it through a dense layer, and then re-adds that value to the output, thus maintaining the original data distribution.
Although the NLinear structure is simple and contains only one layer, it has been shown by several studies to outperform more complicated transformer-family models designed for multi-step time series prediction, such as Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), and FedFormer (Zhou et al., 2022).
Figure 7: The architecture of NLinear model.
@@ 3   Results and Discussion
3.1 Model Configuration and Experimental Design
To validate the efficacy of the SPINN framework, a series of experiments are proposed in this section. The detailed model configuration and comparison experiments are introduced as follows.
3.1.1 Model Configuration and Training Data
Based on the top model configurations identified through comprehensive hyperparameter tuning performed in Hashem and Yousefpour (2024), the $M_{in}$ and $M_{out}$ for all models are set to 168 timesteps, equivalent to seven days. This means that after training the models with the historical time series data up to the current time, the last one-week data is used as input to predict the scour depth for the coming week. Fig. 8 shows the historical time series data of Bridge 539 as an example. Multiple data gaps are found within the data, which is mainly due to sensor malfunctions or seasonal freezing in Alaska.
We use a sliding window with a length equal to $M_{in}$ plus $M_{out}$ to slice the data into the sequences for training. The range of available time series data and the number of available sequences over training and test datasets in each bridge is provided in Table 3.
@@ 3.1.1 Model Configuration and Training Data
For each bridge, we have multiple time series data across years. Once the input and output sequence lengths are determined, we create input-output pairs by sliding a window over the continuous time series data, resulting in a set of available samples. We arrange these samples in temporal order and allocate the last 20% for the test set. The first 80% of the samples are then randomly split into training and validation sets in a 3:1 ratio (60% for training and 20% for validation). The model is validated after each training epoch to monitor its performance on unseen data and prevent overfitting. Fig. 8 illustrates the data split for Bridge 539. Similar illustrations for other bridges can be found in the Appendix Section (Figs. A1 to A3).
3.1.2 Experimental Design
A number of experiments were performed to evaluate the performance of the three SPINN variants and their counterpart pure NN (base data-driven) models. The experiments can be categorized into general and site/bridge-specific as shown in Table 4.
Experiments Training Dataset Base Data-Driven Model $\hat{Y}_{s_{PHY}}$
Pure NN (Site-Specific) 212
527
539
742
LSTM
CNN
NLinear
N/A
Pure NN (General) All LSTM
CNN
NLinear
N/A
SPINN-HEC18 (Site-Specific) 212
527
539
742
LSTM
CNN
NLinear
SPINN-TD (Site-Specific) 212
527
539
742
LSTM
CNN
NLinear
SPINN-GTD (General) All LSTM
CNN
NLinear
$\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}p_{1}\left(y_{1_{max}}\right)^{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\alpha}\left(\frac{q_{max}}{p_{2}}\right)^{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\beta}\left(1-e^{-}\frac{t}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}T_{L}}\right)$
The models trained using bridge-specific data will be evaluated with the corresponding test set, while the model trained on all bridges will be evaluated individually for each of the four bridges. Considering the stochastic nature inherent in NN modelling, each experiment is trained and evaluated five times to ensure the robustness of results. The averaged MSE over the validation and test datasets are used to evaluate the performance of models.
In addition to MSE, we also use Mean Absolute Percentage Error (MAPE) and Root Mean Squared Error (RMSE) as performance metrics to evaluate the models. MAPE and RMSE are defined as:
$RMSE=\sqrt{\frac{1}{M_{out}}\sum_{i=0}^{M_{out}}(y_{i}-\hat{y}_{i})^{2}}$ (11)
3.2 Performance Assessment
@@
Fig. 9 presents the performance distribution for all models over the test dataset. The detailed results of experiments are listed in Table LABEL:tab:table_overall_performance. The aggregated MSE of five runs are listed as mean and standard deviation values for each model. In the following subsections, we will discuss the main findings of these experiments.
Figure 9: Comprehensive analysis of model performance across four bridges, showing the test MSE ($m^{2}$) distribution for pure NN models and SPINN variations. The right-top in each subplot shows the zoom-in area of the NLinear group due to their smaller scales.
3.2.1 Site-Specific and General Pure Data-Driven Models
Analysing the predictive performance across bridges among site-specific and general models is crucial before investigating the impacts of NN algorithms. As observed in Fig. 8, and also Fig. A1 to A3, there are significant fluctuations in the bed elevation $E_{bed}$ across the bridges. For instance, Bridge 539 experienced several significant scour events across the years, with some scour depths reaching over five meters. In contrast, the $E_{bed}$ changes at Bridge 527 were relatively modest, with variations consistently below two meters. Such differences are caused by distinct riverbed material geology, geomorphology, river flow characteristics, and bridge design. All of these various factors can have an impact on the scouring processes and result in predictive complexities.
@@ 3.2.1 Site-Specific and General Pure Data-Driven Models
Pure-LSTM Pure-CNN Pure-NLinear
Elev. Site-Specific General Site-Specific General Site-Specific General
212 0.644 0.533/0.069 0.089/0.034 0.142/0.041 0.101/0.033 0.010/0.014 0.013/0.016
527 2.030 0.141/0.010 0.113/0.010 0.055/0.006 0.123/0.010 0.019/0.004 0.049/0.006
539 4.603 0.478/0.537 0.332/0.546 0.283/0.406 0.276/0.485 0.105/0.250 0.168/0.358
742 3.191 0.884/0.127 0.325/0.111 0.214/0.079 0.310/0.106 0.199/0.076 0.310/0.102
The performance measures presented in Table 5 and Fig. 9 reveal that the relative performance of different NN models varies across bridges. MAPE and MSE criteria indicate different orders of performance for the bridges. Based on the average MAPE values across all models, the order of bridges from lowest to highest MAPE is 527, 212, 742, and 539. In contrast, the order of bridges based on average MSE values from lowest to highest is 212, 527, 539, and 742. These different orderings highlight the importance of considering both MAPE and MSE when evaluating model performance across different bridges, as they capture different aspects of prediction accuracy. MAPE is a percentage-based metric that provides insights into the relative magnitude of errors and is particularly useful when comparing models across bridges with varying scales of bed elevations. On the other hand, MSE concentrates larger errors by squaring them, making it more sensitive to outliers. By considering both MAPE and MSE, we can gain a more comprehensive understanding of the model performance. The trend in model performance does not show a correlation with the amount of training data available for each bridge (see Table 3) and can be pertained to various complex factors such as quality of data, scale of variation, amount of missing data, and consistency in seasonal and temporal patterns across years. Nevertheless, these experiments indicate that the bridges with smaller range/scale of variation and fluctuations of bed elevation seem to achieve better prediction performance overall with both MAPE and MSE criteria.
For the general data-driven models (Pure NN-General), we used the collective dataset from all bridges for their training and then evaluated the model for each bridge scour forecasting individually. The interesting outcome is that the general models yielded better and more stable performance (lower MSE and lower standard deviation) compared to site-specific for the majority of models, except for CNN in the case of Bridge 212 and 539 (see Table 5 and Fig. 9). For instance, training the LSTM model with the combined datasets from all bridges led to 63% and 83% reduction in MSE for bridges 742 and 212, respectively, compared to site-specific LSTM models trained with local data only. These two bridges had specifically less amount of site-specific data, which indicates how a larger dataset, even across different bridges, can enhance the generalization of pure NN models. This outcome highlights the potential for developing accurate global scour forecast models applicable to a cluster of bridges in a region.
While general models in this case study proved superior to bridge/site-specific models, the unique scour characteristics of other bridges may still demand the development of site-specific models to achieve optimal results.
3.2.2 SPINNs and Choice of Physics-based Empirical Equation
The integration of physics-based loss components into NN models, including HEC18, TD, and GTD equations, has led to improvement in predictive accuracy in most of the bridges as observed from Fig. 9 and Table 6. Considering all bridges and base models (represented by the five colours within each group in the figure), the SPINN-based methods consistently achieve the lowest MSE value within their respective groups in most cases, demonstrating their superior performance compared to the pure NN models. For instance, all three SPINN variants have achieved lower MSEs in the LSTM group of Bridge 212. Furthermore, in the LSTM group of Bridge 539, the SPINN-TD helped to reduce the MSE from 0.478 to 0.236 (50.63% error reduction). The reference level ($E_{ref}$) for scour depth calculations differs among SPINN variants: SPINN-HEC18 uses the constant as-built bed elevation, while SPINN-TD and SPINN-GTD employ a dynamic approach, updating $E_{ref}$ based on the bed elevation of the first timestep in each input-output sequence pair.
@@ 3.2.2 SPINNs and Choice of Physics-based Empirical Equation
Table 6: Performance of SPINN variants across different bridges, based on MSE ($m^{2}$) values. Elev. represents the maximum bed elevation variation ($m$) over the test dataset.
@@
Elev. LSTM CNN NLinear LSTM CNN NLinear LSTM CNN NLinear
212 0.644 0.443 0.150 0.010 0.136 0.173 0.010 0.144 0.089 0.013
527 2.030 0.135 0.040 0.019 0.084 0.050 0.023 0.122 0.130 0.049
539 4.603 0.388 0.285 0.105 0.236 0.236 0.114 0.325 0.250 0.169
742 3.191 0.388 0.239 0.200 0.260 0.233 0.229 0.346 0.335 0.310
Table 6 provides a comprehensive comparison of the performance of different SPINN variants across the case study bridges. The SPINN-TD model consistently outperforms the other SPINN variants, achieving the lowest MSE for most of the bridges when paired with LSTM and CNN base models. This suggests that incorporating the proposed site-specific time-dependent equation (TD) as a physics-based loss function is more effective compared to the site-specific HEC18 and the generalized version of the TD equation (GTD). However, the degree of performance improvement varies depending on the specific bridge and base NN model. For instance, while SPINN-TD shows significant error reduction for Bridge 539 when paired with LSTM (50.63% reduction) and CNN (16.61% reduction), the improvement is less pronounced for Bridge 527 with the same base models (40.43% and 9.09% reduction, respectively). This variability in performance gain suggests that the effectiveness of the physics-based loss functions is influenced by the inherent characteristics of each bridge dataset.
Among the different base DL models, LSTM, followed by CNN, seems to achieve the most performance improvement in terms of accuracy (error) and robustness (variance) from a physics-informed architecture. It is noted that the NLinear model does not show a significant benefit from adding the physics loss to its architecture in any form.
3.2.3 Effect of Base DL Algorithm on SPINNs
This paper employs LSTM, CNN, and NLinear as the base models. A comparative analysis of their trainable parameters and computational intensity, expressed in terms of Floating Point Operations (FLOPs), is presented in Table 7. The number of trainable parameters reflects the model’s capacity to learn from data, while FLOPs provide insight into the computational demands of processing that data. In deep learning, FLOPs are estimated by counting the total number of multiplication and addition operations required for one forward pass through the model.

### v3

@@ 2.3.4 Base DL Algorithm: 1-Long Short-Term Memory Networks (LSTM)
In this study, we implement a one-layer LSTM model with 128 hidden units as shown in Fig. 5. The input sequence is processed by the LSTM layer, and then passed through a dense (fully connected) layer with a linear activation function and finally reshaped into the output sequence.
Readers are referred to Yousefpour and Correa (2023) and Hashem and Yousefpour (2024) for more details on optimisation and hyperparameter tuning for the LSTM scour models.
Figure 5: The architecture of LSTM base models for SPINNs, showing the LSTM Memory Unit gates and the input ($x_{t}$), hidden ($h_{t}$), and cell state ($c_{t}$) vectors.
2.3.5 Base DL Algorithm: 2-Convolutional Neural Network (CNN)
@@
In this study, we develop a CNN model with two 2D convolutional layers to show the adaptability of CNNs for processing multidimensional time series data. The two CNN layers have a kernel size of 5, padding size of 2, and output sizes of 128 and 256, respectively. Also, batch normalization layers are added after these layers, and one following dense layer is responsible for producing the final prediction. The CNN model architecture is shown in Fig.  6.
Readers are referred to Hashem and Yousefpour (2024) for details of model optimisation and hyperparameter tuning for the CNN models.
Figure 6: The architecture of the base CNN models for SPINNs.
@@ 3   Results and Discussion
3.1 Experimental Design
A number of experiments were performed to evaluate the performance of the three SPINN variants and their counterpart pure NN (base data-driven) models. The experiments can be categorized into general/transferable and site/bridge-specific as shown in Table 3.
Experiments Training Dataset Base Data-Driven Model $\hat{Y}_{s_{PHY}}$
Pure NN (Site-Specific) 212
527
539
742
LSTM
CNN
N/A
Pure NN (General) All LSTM
CNN
N/A
SPINN-HEC18 (Site-Specific) 212
527
539
742
LSTM
CNN
SPINN-TD (Site-Specific) 212
527
539
742
LSTM
CNN
SPINN-GTD (General) All LSTM
CNN
$\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}p_{1}\left(y_{1_{max}}\right)^{-\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\alpha}\left(q_{max}\right)^{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\beta}\left(1-e^{-}\frac{t}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}T_{L}}\right)$
The models trained using bridge-specific data will be evaluated with the corresponding test set, while the model trained on all bridges will be evaluated individually for each of the four bridges. Considering the stochastic nature inherent in NN modelling, each experiment is trained and evaluated five times to ensure the robustness of the results. The averaged MSE values over the validation and test datasets are used to evaluate the performance of models.
In addition to MSE, we also use Mean Absolute Percentage Error (MAPE) and Root Mean Squared Error (RMSE) as performance metrics for evaluating the models. MAPE and RMSE are defined as:
$RMSE=\sqrt{\frac{1}{M_{out}}\sum_{i=0}^{M_{out}}(y_{i}-\hat{y}_{i})^{2}}$ (11)
3.2 DL Model Configuration and Training
Fig. 7 shows the historical time series data of Bridge 539 as an example (see Appendix Section for time series data for other case study bridges). Multiple data gaps are found within the data, which is mainly seasonal freezing in Alaska. We use a sliding window with a length equal to $M_{in}$ plus $M_{out}$ to slice the data into the sequences for training.
Based on the top model configurations identified through comprehensive hyperparameter tuning performed by Yousefpour and Correa (2023) and Hashem and Yousefpour (2024), the input width, $M_{in}$ and output/label width, $M_{out}$ for all models are set to 168 hours (seven days). This means that after training the models with the historical time series data up to the current time, the last one-week data is used as input to predict the scour depth for the coming week.
@@ 3.2 DL Model Configuration and Training
For each bridge, we have time series data across several years. We create input-output pairs by sliding a window over the continuous time series data, resulting in a number of data slices. We arrange these slices in temporal order and allocate the last 20% for the test set. The first 80% of the samples are then randomly split into training and validation sets in a 3:1 ratio (60% for training and 20% for validation). The model is validated after each training epoch to monitor its performance on unseen data and prevent overfitting. The range of available time series data and the number of available sequences over training/validation and test datasets in each bridge is provided in Table 4. Fig. 7 illustrates the data split for Bridge 539. Similar illustrations for other bridges can be found in the Appendix Section (Figs. A1 to A3).
3.3 DL Model Performance Variation Across Case Studies
@@
Fig. 8 presents the boxplots of performance (MSE) for all models over the test dataset. The detailed performance of all variants is provided in Table LABEL:tab:table_overall_performance. The aggregated MSE of five runs is listed as the mean and standard deviation values for each model.
Table 5 presents the MAPE and MSE criteria for pure data-driven models across case study bridges. It is shown that the order of performance across four bridges is different for the two performance criteria. MAPE is a percentage-based metric that provides insights into the relative magnitude of errors and is particularly useful when comparing models across bridges with varying scales of bed elevations. On the other hand, MSE concentrates larger errors by squaring them, making it more sensitive to outliers. Also, the trend in model performance across different bridges does not show a correlation with the amount of training data available for each bridge (see Table 4).
The DL model performance for each bridge can be influenced by data quality, scale of variation, amount of missing data, and consistency in seasonal and temporal patterns across years. As observed in Fig. 7, and also Fig. A1 to A3, the degree of fluctuations in the bed elevation, $E_{bed}$, varies across case study bridges. For instance, Bridge 539 experienced several significant scour events across the years, with some scour depths reaching over five meters. In contrast, bed elevation changes at Bridge 527 were small, with variations consistently below two meters. Similarly, bridge 212 does not show significant scour and filling throughout the years. It should be noted that this variability in scale and rate of variation in scour depth can affect the relative performance of different DL algorithms. This complexity should be acknowledged in comparing and assessment of competing models.
Figure 8: Model performance across four case study bridges, showing the test MSE ($m^{2}$) boxplots for site-specific (blue) and general (green) Pure NN and SPINN variants.
@@
Pure-LSTM Pure-CNN
Elev. Site-Specific General Site-Specific General
212 0.644 0.533/0.069 0.089/0.034 0.142/0.041 0.101/0.033
527 2.030 0.141/0.010 0.113/0.010 0.055/0.006 0.123/0.010
539 4.603 0.478/0.537 0.332/0.546 0.283/0.406 0.276/0.485
742 3.191 0.884/0.127 0.325/0.111 0.214/0.079 0.310/0.106
3.4 Assessment of SPINN Variants
SPINNs show improvement in predictive accuracy compared with pure NN models in majority of cases. As observed from Fig. 8, SPINNs help with reducing mean error and/or reducing the standard deviation of error for 9 out of 16 experiments (4 bridges, 2 base models, site-specific and general). Refer to Table LABEL:tab:table_overall_performance for a summary of performance (MSE) from all cases/experiments. Table 6 provides mean and standard deviation of MSE for SPINN variants versus their corresponding pure NN models for bridge 539 and 742.
[…]
For the general/transferable models, SPINN-GTD shows comparable performance with marginal improvement. The best outcome is observed for the CNN model of bridge 539 with 9% (from 0.276 to 0.250) MSE reduction.
Table 6: Performance of SPINN variants and Pure NN models, showing MSE ($m^{2}$) in bed elevation prediction over test datasets for Bridges 539 and 742.
Model Training Datasets Method MSE - 539 MSE - 742
LSTM Site Specific Pure NN 0.478±0.157 0.884±0.589
LSTM Site Specific SPINN-HEC18 0.388±0.078 0.388±0.036
LSTM Site Specific SPINN-TD 0.236±0.026 0.260±0.021
LSTM General Pure NN 0.332±0.031 0.325±0.021
LSTM General SPINN-GTD 0.325±0.037 0.346±0.045
CNN Site Specific Pure NN 0.283±0.069 0.214±0.037
CNN Site Specific SPINN-HEC18 0.285±0.070 0.239±0.055
CNN Site Specific SPINN-TD 0.236±0.015 0.233±0.033
CNN General Pure NN 0.276±0.022 0.310±0.040
CNN General SPINN-GTD 0.250±0.022 0.335±0.024
@@ 3.4 Assessment of SPINN Variants
Table 7: Comparison of SPINN variants across four case study bridges, based on MSE ($m^{2}$) values. Elev. represents the maximum bed elevation variation ($m$) over the test dataset.
@@
Elev. LSTM CNN LSTM CNN LSTM CNN
212 0.644 0.443 0.150 0.136 0.173 0.144 0.089
527 2.030 0.135 0.040 0.084 0.050 0.122 0.130
539 4.603 0.388 0.285 0.236 0.236 0.325 0.250
742 3.191 0.388 0.239 0.260 0.233 0.346 0.335
Table 7 compares the performance of different SPINN variants across the four case study bridges. The SPINN-TD model consistently outperforms the other SPINN variants, achieving the lowest MSE for most of the bridges when paired with LSTM and CNN base models. This suggests that incorporating the proposed site-specific time-dependent equation (TD) as a physics-based loss function is more effective compared to the site-specific HEC18 and the generalized version of the TD equation (GTD). However, the degree of performance improvement varies depending on the specific bridge and base NN model. For instance, while SPINN-TD shows significant error reduction for Bridge 539 when paired with LSTM (50% reduction) and CNN (17% reduction), the improvement is less pronounced for Bridge 527 with the same base models (40% and 9% reduction, respectively). This variability in performance gain suggests that the effectiveness of the physics-based loss functions is influenced by the inherent characteristics of each bridge dataset.
3.5 Site-Specific vs General/Transferable DL Models
The general pure data-driven models (Pure NN) show better performance (lower MSE and lower standard deviation) compared to site-specific models, except for CNN in the case of Bridge 212 and 539 (see Table 5 and Fig. 8). For instance, training the LSTM model with the combined datasets from all bridges led to $\sim$60and $\sim$80reduction in MSE for bridges 742 and 212, respectively. It is noted that these two bridges had less amount of site-specific data. This indicates aggregating data from a cluster of similar bridges can potentially enhance the generalization of the DL models. This outcome highlights the potential for developing transferable scour forecast models applicable to a cluster of bridges in a region.
However, the general hybrid models, that is, SPINN-GTD did not show consistent superiority over site-specific hybrid models, i.e. SPINN-HEC18 or SPINN-TD (except for 539 LSTM, 742 LSTM, 212 CNN cases). The challenge of transferability in the SPINN-GTD model is that the general empirical equation is being fit to a wider range of scour characteristics. The GTD equation in fact combines all the site/bridge-specific coefficients and inputs of the site-specific TD equation (i.e., riverbed, flow and pier shape/dimension factors) into one latent parameter ($p_{1}$). This could limit the extent to which the empirical model can help NN learn and adapt adequately to the inherent scour patterns for each bridge, thus reducing the prediction accuracy.
3.6 Evaluation of Base DL Algorithms for SPINNs
As shown in Table 5 the relative performance of pure DL models varies across bridges, however, CNN seems to overall outperform LSTM in prediction accuracy, however not in computational cost as shown in Table 8.
The comparison of computational cost between LSTM and CNN is expressed in terms of Floating Point Operations (FLOPs). In deep learning, FLOPs are estimated by counting the total number of multiplication and addition operations required for one forward pass through the model. The number of trainable parameters reflects the model’s capacity to learn from data, while FLOPs demonstrate the computational demand for training.